OpenStudy (anonymous):

How would is solve 5e^0.2x=7 Does it become 5^0.2x=ln7

OpenStudy (phi):

\[ 5e^{0.2x}=7 \] you could do this: \[ \ln(5e^{0.2x}) = \ln(7) \] but that is not the best thing. There are rules for tackling this , but it's easier if we do this instead: \[ 5e^{0.2x}=7 \] divide both sides by 5 \[ \frac{5}{5} e^{0.2x}= \frac{7}{5} \] 5/5 is 1 so the left side simplifies to \[ e^{0.2x}= \frac{7}{5} \] now write ln around each side